# Illuminating Blind Spots: Exploring LLMs as a Source of Targeted Synthetic Textual Data to Minimize High Confidence Misclassifications

### v2

@@ 3.3 Synthetic Sample Generation
Figure 2: Workflow: (a) Attacking the original model to obtain UUs; (b) Using these UUs to extend the training data via generalization (figure 1) and thus obtain a more robust model; (c) Evaluating this retrained model.
4 Experimental Setup
Initially, we provide an overview of the overall experimental setup, followed by a detailed description of the setup for our experiments, shown schematically in figure 2. Additionally to addressing natural blind spots (section 5), we performed a study on human-induced blind spots (appendix A), in which we show that one can create blind spots in a model that do not naturally occur when training on the data set, and subsequently successfully target its UUs via our approach, thus eliminating that blind spot. For further contextualization of the data we collect, we show a selection of data for the IMDB task. In table 3, we present an example of hypothesis generalization from a UU via abstraction and in table 4 an example of hypothesis generalization from a previous hypothesis by exploration.
@@ 4 Experimental Setup
Table 2: Datasets used, including the task type, number of classes and number of samples in each of the test, validation, and training sets. Note the split of the original IMDB test set into validation and test set.
Original/Perturbed Samples O: […] Footage of busy highways, planes flying overhead, poor lightning effects, soldiers doing their thing, they’re all used in a [[haphazard]] manner, sometimes repeatedly, and they often only have a loose connection to the story Wood is trying to tell. […] P: […] Footage of busy highways, planes flying overhead, poor lightning effects, soldiers doing their thing, they’re all used in a [[thoughtless]] manner, sometimes repeatedly, and they often only have a loose connection to the story Wood is trying to tell. […]
Human Abstraction Hypothesis “Haphazard can be intentional, vs thoughtless is unplanned. It feels more of an unwillingness rather than a lack of ability”
LLM Abstraction Hypothesis “Hypothesis: The replacement of specific adjectives in the perturbed sample alters the sentiment and contributes to misclassification. Reasoning: The adjective "thoughtless" in the perturbed sample, replacing "haphazard" may change the perception of the director’s skills and use of stock footage. This alteration in sentiment could lead to a different classification outcome.”
Table 3: Examples of hypothesis generalization using abstraction for the IMDB dataset. Starting from a UU made up of an original and its corresponding perturbed sample, abstraction is performed by a human or LLM.
Human Abstraction Hypothesis “The pattern that may have caused the wrong prediction may be related to the substitution of the words with their synonym, for example ""bad"" with it’s synonym ""wicked"".”
Human Exploration Hypothesis “The hypothesis could be that the words were potentially swappped or replaced, such as ""bad"" been replaced by ""evil"" which could lead to an incorrect prediction.”
LLM Abstraction Hypothesis “The change in the word "appalling" to "horrifying" may have caused the misclassification. Both "appalling" and "horrifying" have negative connotations and can be used to describe something that is disturbing or shocking. However, the word "appalling" may be considered more extreme and may have a stronger negative connotation compared to "horrifying". This subtle difference in the intensity of the negative sentiment may have caused the misclassification by the algorithm.”
LLM Exploration Hypothesis “The use of the word "hypocrite" may have caused the misclassification. Reasoning: The word "hypocrite" implies that the person is saying one thing but doing another, which may be considered a negative trait. However, some people may not interpret the person in the sample as a hypocrite, leading to a difference in sentiment analysis. This difference in interpretation may have caused the misclassification by the algorithm.”
Table 4: Examples of hypothesis generalization using exploration for the IMDB dataset. Starting from a hypothesis obtained via abstraction, exploration is performed by a human or LLM.
4.1 Verification
For quality control purposes of the human contributions we include two attention checks in each survey to determine if crowdworkers are paying attention Oppenheimer et al. (2009). Additionally, we eliminate low-effort responses submitted by crowdworkers. Our goal for this is not to select high-quality responses, but instead to eliminate answers that were made in bad faith, e.g. by copy-pasting the same response repeatedly. For automated quality control for responses by humans and LLMs our approach is two-fold. For all generated text, whether hypothesis or sample, we ensure a certain character length ($\text{char}_{\text{min}}=40$) to exclude responses that have insufficient usable content. Finally, we employ an automatic quantitative evaluation technique in the form of BERTScore (Zhang et al., 2020). Here, we use the handcrafted hypothesis set and generations from appendix A as the reference set. If the score of a new hypothesis or sample fall below a threshold $S_{min}$, then we exclude it and repeat the previous step to obtain a replacement. In practice, we find that $S_{min}=0.5$ gives a good balance between accuracy and ensuring textual diversity.
4.2 Datasets and Models
We systematically evaluate the effectiveness of our approach on a range of distinct classification tasks, which were selected gauge the adaptability of our approach across different levels of task complexity. We focus on sentiment analysis (SA), semantic equivalence (SE), and natural language inference (NLI) for our experimental evaluation. To this end, we use the popular IMDB Maas et al. (2011) dataset for the SA task, the MRPC Dolan and Brockett (2005) dataset for the SE task, and the QNLI Rajpurkar et al. (2016) datasets for the NLI portion of our experiment. An overview of the details of the used datasets can be found in table 2. We choose this combination of datasets because they exhibit significant disparities in the training sample sizes, along with varying levels of task complexity. We finetune a BERT Devlin et al. (2019) model to perform the classification task and use the bert-base-uncased implementation provided by Wolf et al. (2020).
For blind spot mitigation, we restrict the number of hypotheses to 1% of the training set size for both abstraction-derived and exploration-derived hypotheses. Additionally, each generated sample corresponds to a single hypothesis, and we utilize each hypothesis, resulting in the number of new samples being equal to 2% of the training set size. These values can be effectively regarded as budget-dependent hyperparameters. The even split between abstraction-derived and exploration-derived hypotheses is also a choice that can be optimized, which we leave for future work. In line with previous works Lakkaraju et al. (2017); Bansal and Weld (2018); Han et al. (2021), we define $\tau=0.65$. For the generation of hypotheses or samples via LLM, we use GPT version gpt-3.5-turbo-1106 at a temperature $T=0.7$ with the default system prompt. All models are trained using a learning rate of $2\times 10^{-4}$ and a batch size of 64 for 10 epochs.
4.3 Perturbations
We assume no knowledge of the model parameters, i.e. a black-box setting, and focus on perturbation techniques that are most likely to maintain the same meaning post perturbation and thus preserve the underlying true label. We implement two state-of-the-art perturbation techniques via the TextAttack framework Morris et al. (2020), namely TextFooler Jin et al. (2020) and DeepWordBug Gao et al. (2018). TextFooler focuses on semantically and syntactically similar synonym swap perturbations. It identifies the most important words for the target model and then replaces them with semantically similar and grammatically correct words until the prediction is altered. DeepWordBug instead focuses on a range of character level perturbations that aim to minimize the edit distance of the perturbation. We thus cover two major lines of adversarial perturbations, i.e. world-level and character-level. Note that perturbation methods are not a necessity to uncover blind spots. Instead, we employ them to perform proactive instead of reactive illumination of blind spots and thus find additional instances failure modes.
4.4 Evaluation
We explore a range of metrics to evaluate the effectiveness of our approach. Specifically, we assess the accuracy of a model, the percentage of successful perturbations achieved by the perturbation technique, and the number of UUs that occur due to the perturbations. All of these are recorded for the model trained on the original dataset and for the one retrained on $\mathcal{E}$. The overall accuracy of the model on the test set gives basic insight into the general performance, while the successful perturbation percentage and the number of UUs indicate the overall robustness of the model and prevalence of blind spots that lead to high confidence misclassifications, respectively. It is desirable to maximize the first while minimizing the latter two. As an increasing number of blind spots are illuminated in the classification model, the number of UUs is reduced. As such, the model is less prone to catastrophic misclassifications due to perturbation attacks or other out-of-distribution data.
5 Natural Blind Spots
Unlike synthetic blind spots, which are induced via biasing, natural blind spots are present in the model due to natural biases in the train set. Beyond studying whether our approach is indeed successful in reducing blind spots across our diverse choice of classification tasks, we aim to compare human to LLM performance in terms of effectiveness, ability to scale, and ease of use.
One crucial difference in instruction between the human and LLM procedures is the number of examples given, with humans receiving two, fully worked out examples per participant, while no example is given to the LLM. This was done to lower the amount of guidance given to the model, as few-shot prompting the LLM leads to a highly uniform set of samples, even at higher temperature values. We thus intentionally introduce more variability, akin to what naturally transpires in the human procedure, where crowdworkers possess distinct perspectives and thoughts. The structure and content of the user study presented to the crowd workers is shown in appendix C, while the prompt structure used for the LLM is shown in appendix D.
TextFooler DeepWordBug Acc. (%) $\uparrow$ UUs (#) $\downarrow$ Acc. (%) $\uparrow$ UUs (#) $\downarrow$ MRPC Original 82.38 952 82.38 936 LLM Data 81.57 851 82.23 882 Human Data 81.58 418 82.10 802 IMDB Original 94.84 1882 95.40 1682 LLM Data 95.40 1241 94.41 1448 Human Data 94.43 1518 95.74 1412 QNLI Original 89.88 1923 89.88 2597 LLM Data 89.31 1536 89.38 1746 Human Data 89.42 2028 89.38 2325
Table 5: Results of natural blind spot study for accuracy, perturbation success rate, and number of UUs. Rows show results for the original and retrained models (using human- or LLM-generated synthetic samples) by dataset.
5.1 Natural Blind Spot Study Results
The results for the addition of the human- and LLM-generated samples are shown in table 5.
Accuracy and UUs Regarding accuracy, which serves as a proxy for general performance, we find that the procedure employed, whether human- or LLM-based, has no significant impact. Across the various datasets, the accuracy remains consistent, with a difference of no more than $\pm$ 1% between the original and retrained models. With regards to our main metric, the number of UUs, we observe that our method can successfully lead to decreases, with a maximum reduction of 56.09% for the human-based retraining on the MRPC dataset using TextFooler. The only exception to this is the human-based procedure on the QNLI dataset with the TextFooler attack, where we saw an increase of 5.46%. On average, across the two perturbation methods, there is a reduction in UUs of 19.54% and 16.80% for LLM-based and human-based retraining, respectively.
Perturbations A detailed overview of the perturbation statistics can be seen in appendix B. Perturbation success rate, which serves as a proxy for general model robustness against attacks, indicates that leveraging the LLM-based method leads to more stable results, with reductions of up to 8.82% for the IMDB dataset using DeepWordBug, while the human-based data leads to wider swings in performance.
Human Versus LLM Performance The decrease in UUs for LLMs is more substantial compared to the decrease observed as a result of the use of human-generated samples, as well as more predictable, suggesting that LLMs create quality hypotheses and samples more consistently. However, in the case of the TextFooler perturbation method on the MRPC dataset, the human-based approach exhibits a remarkable reduction in UUs, outperforming the LLM-based one significantly. This indicates a strong task dependency of the effectiveness of humans in characterizing blind spots using our proposed approach.
The task-dependent variability in human performance is evident in the undesired outcomes for the QNLI dataset when using TextFooler. Contrary to expectations, the responses generated by humans actually result in an increase in UUs. The observed variability in performance can likely be attributed to the intuitiveness of the SE and SA tasks, which contrasts with the complex NLI task. Determining the equivalence of two sentences or discerning the sentiment of a movie review aligns with the everyday experience of individuals. As a result, these tasks are perceived as less complex and more intuitive to crowdworkers. This underscores the task dependency of non-expert human abilities and highlights how human performance can vary based on the nature of the task at hand.
5.2 Distribution of Successful Perturbations
To visualize the effect of the retraining across the perturbed samples, we track the confidence at which predictions are made. The confidence plots across datasets are shown in figure 3 and figure 4 for the TextFooler and DeepWordBug perturbation methods, respectively. We can observe a reduction in high confidence predictions across the retrained models; specifically, predictions with confidence of above 90% are reduced as a result of retraining. Additionally, we observe similar results across perturbation methods, indicating that the performance of our approach is not dependent on the type of perturbation method employed.
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100200200200200400400400400Confidence# of Instances
(a) MRPC Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100200200200200400400400400Confidence
(b) MRPC Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100200200200200400400400400Confidence
(c) MRPC Retrain Human
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Confidence# of Instances
(d) IMDB Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Confidence
(e) IDMB Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Confidence
(f) IMDB Retrain Human

### v3

@@ 2.3. Synthetic Sample Generation
Figure 3. Workflow: (A) Obtain UUs from the validation set on the original finetuned model; (B) use UUs to extend the training data via generalization (figure 1) and thus obtain a more robust model; (C) evaluate this retrained model. Adversarial perturbations in dotted box are optional.
3. Experimental Setup
In this section, we present an overview of our experimental design. A schematic illustration of the workflow can be found in figure 3. Here we first obtain our initial set of UUs of the finetuned classification model from the validation set. Following this, we characterize the blind spots corresponding to these UUs by making the intelligent agent perform generalization as described in section 2, culminating in new synthetic data that we use to retrain the model. Finally, we evaluate this retrained model with respect to accuracy and UU count. As a preliminary study, to verify that our method does indeed address blind spots, we successfully demonstrate that it is possible to artificially create blind spots by hand (i.e., ground truth blind spots) in a model and then illuminate these using our approach in appendix A. In our main study, our experiments instead address mitigating both natural blind spots that occur during normal model training and those created by adversarial attacks. For this, we do not have access to the ground truth blind spots and as such just have indirect evidence that some blind spots are illuminated as the number of occurring UUs is decreased.
3.1. Datasets, Models, and Perturbations
To evaluate the generality and effectiveness of our approach, we select a diverse set of classification tasks, each representing varying levels of task complexity. Specifically, we focus on sentiment analysis (SA) using the IMDB dataset (Maas et al., 2011), semantic equivalence (SE) using the MRPC dataset (Dolan and Brockett, 2005), and natural language inference (NLI) using the QNLI dataset (Rajpurkar et al., 2016). The statistics of the dataset for each task are shown in table 1. For blind spot mitigation, we use the validation set to obtain our UUs that are then used to perform the hypotheses generalization. These hypotheses are then used in turn to generate synthetic samples and extend the training set, as shown in figure 3. We limit the number of hypotheses derived from each of abstraction and extrapolation to 1% of the training set size, leading to an additional 73, 500, and 2095 training samples after applying our method for MRPC, IMDB, and QNLI, respectively. These values are treated as hyperparameters and are chosen to balance computational efficiency and effectiveness. We leave further optimization of this split between abstraction- and extrapolation-derived hypotheses to future work. We employ two classification models in our experiments, finetuned for each classification task: BERT (bert-base-uncased(Devlin et al., 2019) and Llama 2 (llama-2-7b(Touvron et al., 2023), selected for their contrasting architecture and size. We choose BERT for its known performance on sentence-level tasks and its low number of parameters, while Llama 2 was chosen for its larger (but still manageable) scale and capability in handling more complex language understanding tasks. GPT-3.5 (gpt-3.5-turbo-1106(Brown et al., 2020b) is incorporated as the teacher model to perform hypothesis and sample generation, as it is superior to both classification models that we use.
In a black-box setting, where we assume no access to the model’s internal parameters, we employ adversarial perturbation techniques to yield more UUs for our method to use. Note that while perturbations aid proactive discovery of blind spots, they are not strictly necessary to our overall approach. Perturbations are generated using TextAttack (Morris et al., 2020), specifically with TextFooler (TF) (Jin et al., 2020) for word-level perturbations and DeepWordBug (DWB) (Gao et al., 2018) for character-level perturbations. Using these two methods, we cover a wide spectrum of adversarial attack types, revealing additional blind spots. We focus on perturbations that maintain semantic integrity, ensuring that the true underlying label remains consistent after perturbation. Manual inspection of 100 random perturbed samples revealed that none had a different underlying true label, affirming that our perturbations are faithful.
3.2. Baseline
As a baseline, we use a reactive relabeling approach based on the previous work by  (Han et al., 2021), where identified UUs are given a ground truth label, before being reintroduced to the classification model for additional training. This method directly targets blind spots by adding these correctly labeled samples to the extended set. While  (Han et al., 2021) performs this reintroduction in smaller, iterative batches to identify more UUs, we pool all relabeled UUs in a single batch, as we only concern ourselves with the mitigation of UUs and assume that we have knowledge of whether a sample is a UU or not post classification. This is similar to how we perform the retraining for our method. For a fair comparison, we apply this baseline approach with the same budgetary constraints as our proposed method, with new samples making up 2% of the initial training set size. We pose that our method, which uses hypotheses to synthesize new data, will outperform this method by uncovering additional failure modes not captured by relabeling alone.
3.3. Implementation
Following Lakkaraju et al. (2017), we set the confidence threshold for determining high-confidence misclassifications to $\tau=0.65$. We use GPT-3.5 with a temperature setting of $T=0.7$ and the default system message. All BERT models were trained for 10 epochs, using a learning rate of $2\times 10^{-4}$, and a batch size of 64. We fine-tune all Llama 2 7B models using the Low-Rank Adaptation (LoRA) (Hu et al., 2021) method with the following configuration: a LoRA scaling factor of 16, dropout of 0.1, and rank $r=64$. The target modules are all linear layers in the model, and no bias adjustment is applied. The training is performed over 3 epochs, with a batch size of 8, and gradient accumulation set to 8 steps. We employ AdamW as optimizer. The learning rate is set to $2\times 10^{-4}$ with a cosine learning rate schedule and a warmup ratio of 0.03. We apply a maximum gradient norm of 0.3 to ensure stability during training. We use a weight decay of 0.001 to prevent overfitting.
The human computation component of our study is implemented through a survey study, the details of which are provided in appendix B. A key procedural difference between human and LM-based experiments is the number of examples provided. The human participants receive two examples, while no examples are given to LMs (i.e., zero-shot). This design choice aims to minimize guidance for the LM since few-shot prompting tends to result in overly homogeneous samples, even when using higher temperature settings. The LM prompts for the teacher model are given in appendix C. When prompting the teacher model, we always ask it to explicitly give its reasoning, which we find not only increases performance but also improves interpretability. To ensure the quality of human-generated hypotheses and synthetic samples, we include attention checks (Oppenheimer et al., 2009) in each survey to eliminate inattentive or low-effort responses. For both human- and LLM-generated hypotheses and samples, we implement automated quality checks for this purpose. We do not focus on selecting the high-quality responses, but filter out bad-faith ones such as repeated or nonsensical submissions. To be included, all text entries are required to meet a minimum character threshold ($\text{char}_{\text{min}}=40$) to ensure sufficient content. Additionally, we employed BERTScore (Zhang et al., 2020) to automatically evaluate the similarity of new samples against a reference set in the form of samples from the training set. If the similarity score falls below a threshold of $S_{min}=0.5$, the entry is discarded.
Table 1. Datasets used, including the task type, number of classes, and number of samples in each of the test, validation, and training sets. Note the split of the original IMDB test set into new validation and test sets.
@@ 3.3. Implementation
3.4. Evaluation Metrics
We use two key metrics to assess the effectiveness of our approach and the comparative approach. These include the accuracy of the model on the test set and the number of UUs observed during evaluation. Accuracy provides a basic measure of model performance, while the UU count reflects the model’s robustness and allows us to reason about the prevalence of blind spots. Note that the accuracy we report is the accuracy of the model before any perturbations are applied, while the number of UUs is post perturbation. Ideally, our goal is to maximize accuracy while minimizing the number of UUs. Our evaluation compares the performance of the original finetuned model with that of the models retrained on their respective extended dataset $\mathcal{E}$. This allows us to quantify the impact of our approach on mitigating blind spots and improving model robustness.
Table 2. Results of the blind spot study across datasets for BERT and Llama 2 7B as classification models. Here TF refers to the TextFooler perturbation method and DWB to DeepWordBug. An $\uparrow$ indicates that a higher score is preferable, while $\downarrow$ indicates that lower is better.
BERT Llama 2 7B
TF DWB TF DWB
Acc (%) $\uparrow$ UUs (#) $\downarrow$ Acc (%) $\uparrow$ UUs (#) $\downarrow$ Acc (%) $\uparrow$ UUs (#) $\downarrow$ Acc (%) $\uparrow$ UUs (#) $\downarrow$
MRPC Original Model 82.38 952 82.38 936 90.84 301 90.66 293
Relabelling Baseline 82.49 911 82.55 898 90.61 277 90.73 268
Our Method w/GPT-3.5 81.57 851 82.23 882 89.86 149 89.73 164
Our Method w/Humans 81.58 418 82.10 802 90.20 144 89.91 140
IMDB Original Model 94.84 1882 95.40 1682 95.20 892 95.33 810
Relabelling Baseline 93.94 1732 94.26 1621 94.86 781 95.10 742
Our Method w/GPT-3.5 95.40 1241 94.41 1448 94.96 604 95.13 689
Our Method w/Humans 94.43 1518 95.74 1412 94.67 658 94.90 702
QNLI Original Model 89.88 1923 89.88 2597 90.08 879 90.72 952
Relabelling Baseline 88.24 1796 88.98 1907 89.90 856 90.60 929
Our Method w/GPT-3.5 89.31 1536 89.21 1746 89.58 741 90.10 890
Our Method w/Humans 89.42 2028 89.38 2325 89.16 857 89.73 924
4. Results
In this section, we report the experimental results on the effectiveness of our proposed method in reducing blind spots across the classification tasks. The results of our methods configured with human- and LM-generated data as well as those of the baselines are shown in table 2. Additionally, we compare human-generated samples to those produced by LMs in terms of effectiveness, scalability, and ease of use.
4.1. Impact of Synthetic Samples
4.1.1. Observation 1: Our approach leads to a significant and consistent UU reduction across tasks
As part of our evaluation, we find that our method successfully reduces UUs, with a maximum reduction of 56.09% when using human computation on the BERT model with TF for the MRPC task. On average, across perturbation methods and classification models, our method with LM-based data generation reduced UUs by 22.37%, while human-based data generation led to a reduction of 15.78%. Similarly, regardless of what type of agent generates the data, our method achieves a reduction in UUs of 35.77%, 21.46%, and 13.03% for MRPC, IMDB, and QNLI, respectively. These results highlight the strengths of using agent-generated samples, with large LMs as a teacher model generally offering more consistent reductions in UUs, though there are difference between tasks. The only configuration where our method does not reduce UUs is the BERT model on the QNLI dataset, where human-based retraining with TF actually increases UUs by 5.46%. We elaborate on this in observation 3.
4.1.2. Observation 2: Relabeling of UU samples is effective but not as impactful
Simply relabeling UU samples from the validation set and reintroducing them as the extended set leads to a decrease in the number of UUs, albeit a more modest one compared to our method. Relabeling achieves a consistent decrease in UUs across tasks of 11.10% and 7.08% on average for BERT and Llama 2, respectively, compared to an average decrease of 14.10% and 17.45% for our method when using humans and 22.74% and 22.00% when using LMs. This confirms that only reactive illumination of blind spots using seen data is less effective than our method, regardless of agent type, as the characterization and subsequent extrapolation we employ results in a more significant reduction in UUs. While the average decrease is lower, the relabeling method is very consistent across tasks, as it is not dependent on an agent grasping the task and delivering high quality data. Additionally, it is very cost effective as no human computation or LM querying is necessary. The obvious limitation of this approach is that it only scales to blind spots that have been discovered and therefore has very little transfer learning potential, as it is unlikely that the found UUs with generalize to unseen UUs.
4.1.3. Observation 3: Human performance is very task dependent
We find that human-generated samples may outperform LMs in tasks that align with human intuition. For tasks such as SE and SA – which are more intuitive to humans compared to NLI, as they more closely resemble everyday tasks – human performance tends to be better, yielding more significant reductions in UUs. In particular, on the MRPC dataset we see a greater reduction in UUs using human-generated data, 35.38% and 52.19% on BERT and Llama 2, respectively, when compared to when using LM-generated hypotheses and samples 8.21% and 47.31%. In less intuitive tasks such as NLI, humans can generate data of poor quality, leading to a reduction in model robustness, which may even result in an increase in UUs. When analyzing participants’ responses for QNLI, we find that several participants did not fully grasp the natural language inference task, which was not the case for SE and SA. Note that these are not purposefully low-effort responses and are therefore not filtered out as described in section 3.3. This shows that irrespective of classification model, there is a task-specific advantage of human computation compared to LM teacher models when there exists a higher degree of familiarity with the task and vice versa. Although LMs provide samples of acceptable quality consistently, rare but high-quality human responses, such as a crowdworker correctly identifying that changing the date “June 15” to “John 15” referenced a Bible verse – an insight that the LM missed – can significantly reduce UUs and thus be more impactful. This suggests that while human-generated responses can have a higher ceiling in certain contexts, LMs deliver more consistent results overall as just a few human participants’ incorrect responses can reduce the effectiveness of our method.
4.1.4. Observation 4: Accuracy does not decrease despite improved robustness
In terms of accuracy, extending the training set with human- or LM-generated data did not have a significant effect. Across tasks, accuracy fluctuations of the models with extended training sets remain within $\pm$1compared to the original models. This contrasts with previous findings that improvements in robustness often come at the expense of accuracy (Tsipras et al., 2019). To illustrate the impact of retraining on accuracy, we visualize prediction confidences across misclassified samples post perturbation for a selected dataset and perturbation method in figure 4. Here, we observe a similar pattern to all other experimental configurations, namely a reduction in high-confidence misclassifications, particularly at the highest prediction confidences. Additionally, there is a clear reduction across the entire confidence range towards lowering the confidence the classifier model has in its misclassifications. This, in combination with our overall results, indicates that we improve the calibration of the classification models. Detailed perturbation statistics, shown in table 7, further demonstrate that the LM-based method provides more stable robustness improvements.
4.2. Scalability and Ease of Use
4.2.1. Observation 5: Our method scales well per sample and by parameter count
Despite only adding a small amount (2% for each task) of synthetic data relative to the total training set size, we achieve significant results in the reduction of UUs. This indicates that our method can scale to large datasets, as only a small number of synthetic samples relative to the total dataset size are required have a significant impact in terms of improving robustness. We study classification models that use a different architecture and have an order of magnitude difference in size (110M parameters for BERT and 7B for Llama 2). Here, we find that models with a lower number of parameters achieve a performance similar to that of large generative LMs, with comparable accuracy on the IMDB and QNLI tasks, indicating that smaller models may be more suitable for text classification tasks when considering their other advantages, which corroborates previous findings (Yu et al., 2023). This is especially encouraging for use cases where computational resources are limited or speed and transparency are critical.
4.2.2. Observation 6: Obtaining samples via LM is easier and more cost effective
When considering the practical aspects of our study, significant insights emerge regarding the costs and time involved in conducting human- and LM-based generalization experiments. The human study, which included 168 participants, resulted in a total cost of \$1072, with an hourly compensation rate of \$12 per participant. In contrast, the LM experiment incurred a much lower cost of \$46 for generating an equivalent number of generalizations and samples. Although it is challenging to provide precise estimates, the data collection process via human surveys also took substantially longer than the LM-based approach. This highlights the fact that when using LMs, our method is far more cost-effective and generates data almost instantaneously, in stark contrast to the considerable delays associated with human-based study design and data collection. Thus, from a scalability perspective, the LM-based procedure offers clear advantages, being both faster and less expensive. However, in certain high-stakes or specialized applications such as suicide prevention and criminal justice sentencing, human involvement, including via a hybrid approach where human intuition supplements the efficiency of LM-generated data, may be more advantageous. This is especially true when considering that LM outputs come with no guarantees and may be biased. These findings underscore the resource implications of choosing between human and LM-based methods, helping researchers plan and allocate resources more effectively.